1.1 Using graphical method, solve the equation; π‘₯

2

- 3π‘₯ – 4 = 0

1.2 Solve the equation; x

2

- 8x + 7 = 0 using the Quadratic Formula


1
Expert's answer
2021-08-23T07:23:08-0400

1.1 Using graphical method, let solve the equation "\ud835\udc65^2- 3\ud835\udc65 -4 = 0."

Let us sketch the graph of the function "f(x)=\ud835\udc65^2- 3\ud835\udc65 -4:"




It follows from the graph that the solutions of the equation "\ud835\udc65^2- 3\ud835\udc65 -4 = 0" are "x=-1" and "x=4."



1.2 Let us solve the equation "x^2- 8x + 7 = 0" using the Quadratic Formula.


Let us find the discriminant:


"D=(-8)^2-4\cdot 7=64-28=36"


We conclude that the solutions are


"x_1=\frac{8+\sqrt{36}}{2}=\frac{8+6}{2}=\frac{14}{2}=7" and "x_2=\frac{8-\sqrt{36}}{2}=\frac{8-6}{2}=\frac{2}{2}=1."